OpenStudy (anonymous):

1. Solve the following equations. Show each step of your work. Check your answer. a) 2x – 3 = -5(x-4) b) 4x2 + 4x = 24

OpenStudy (texaschic101):

2x - 3 = -5(x - 4) --- distribute the -5 through the parenthesis 2x - 3 = -5x + 20 --- add 3 to both sides 2x = -5x + 20 + 3 --- add 5x to both sides 2x + 5x = 20 + 3 -- combine like terms 7x = 23 -- divide both sides by 7 x = 23/7 ======================= 4x^2 + 4x = 24 4x^2 + 4x - 24 = 0 4(x^2 + x - 6) = 0 4(x + 3)(x - 2) = 0 x + 3 = 0 x = -3 x - 2 = 0 x = 2

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (texaschic101):

anytime :)